IPP Progress Report for the week 2010.07.05 - 2010.07.09

(Up to IPP Progress Reports)

Eugene Magnier

With the telescope down this past week due to shutter problems, we have been pushing on other, background processing which is needed. This includes more analysis of the darks (still in progress), working on the deep stack for MD06, and starting the reference stacks for MD08, MD09, and MD10, which are now starting to be observed alot.

I continue to work on extended source analysis. My attempts last week to improve the stability of the fits was not as successful as I hoped. I have turned in another direction: I have modified the analysis for the Sersic model to first do a coarse grid search in the index and normalization, before allowing the rest of the terms to float. I have been very happy with the results of this method so far -- the grid search gets us very close to the required index, and once there the fitting is much more stable. I'll be doing more testing, and concentrating on a number of (unrelated) psphot issues that have been identified before this effort will be finished.

Serge Chastel

  • Still in learning mode
  • Summary plots in ippMonitor:
    • Warp Stage Exposures: Column added
    • Stack Summary w/ Images: New page / .d definition file
    • Diff Summary w/ Images: New page / .d definition file
    • Slight modification to the ippMonitor generation script => images (e.g. part of file name) can now be defined using OP

Heather Flewelling

  • made sky brightness plots for ken (for days since 2010-05-12)
  • detrends
    • queued accidentally a massive number of darks (these were stopped)
    • queued darks with a linear temperature term
    • queued darks with a quadratic temperature term
    • the residuals for both darks look wrong compared to the darks that we are currently using, and I'm currently investigating that.
  • czar for 2 days - nothing exciting to report
  • attended mops meeting
    • processed 2010-05-12 and 2010-05-11 data without dynamic masking
    • learned (from Paul) how to publish & distribute this
  • answered liason questions for LGOCT & IfA

Roy Henderson

  • Short week as I was on vacation on Friday
  • PSPS sandbox testing:
    • Worked with Conrad to get new batches to sandbox with new DXLayer: lots of issues resolved
    • located frames with dodgy astrometry in PSPS and readied for loading to sandbox
  • IPP
    • Czar on Monday, but no data...
    • changes to diff code to get ready to publish to PSPS
    • started work on 'roboczar':
      • created a Db to hold data
      • added polling to my czartool script so that it writes to the Db
      • added plotting function so that time-series plots can be made of any label and stage
      • histograms of exposures processed for each stage in a given time window

Paul Price

  • Stacks:
    • Excess noise in photometry was due to burns (lots in these dithered STS exposures!)
    • Suggest adding BURNTOOL to PPSTACK:MASK.VAL
    • Convolved stacks now have clearly better photometry than unconvolved
    • Made plots and wrote up
  • Noise model: move scaling factors around so variance map can include variation due to kernel variation

Bill Sweeney

  • More debugging of update problems.
  • Modified postage stamp system to detect unrecoverable errors in the update process and fault the jobs. This prevented users from experiencing stuck requests as often.
  • Fixed problem with chip stage reverts causing rows in the database to disappear. Recovered many of the deleted rows from backup database.
  • Started working on building a PS1 reference catalog for MD06 following the procedure that Gene followed for MD04. We have more MD06 images and this has uncovered some problems in the dvo programs that we have to work through.
  • Selected good quality exposures for MD08, MD09, and MD10 in order to build a reference stack for these fields. Queued them for processing which completed through warp, but we decided to redo the tessellation to have overlapping skycells so we need to re-queue the warps.
  • Debugged some problems with detectability requests submitted by MPIA. Passed these along to Chris our expert in this area.

Chris Waters

  • Burntool: wrote task to run burntool on old data using the same syntax as the nightly science processing. Identified bug in burntool queuing, and noted that fpack/funpack is not consistent. Reprocessed two months of data.
  • Fixed bug in summary stage database tables that was preventing a single exposure from creating multiple projection cell images (such as for ThreePi?).
  • Switched detverify to use the original master flats instead of the corrected.
  • Fixed bug in detectability server introduced last time I made a change.
  • Created plots of cell-level linearity study. Identified cells that are not linear.